OpenStudy (chycora):

2. (1 pt) Dora can run 3/5 mi in 3/20 h. How far can she run in 1 h? mi

OpenStudy (chycora):

@Agl202

OpenStudy (agl202):

3/5 mi 3/20 h X mi 1 h ___________ X*3/20 = 1* 3/5 X=4 mi

OpenStudy (chycora):

thanx

OpenStudy (agl202):

:D Welcome!